OpenStudy (anonymous):

Please help! I need to know if I am correct. A lever acts like a fulcrum to help lift heavy objects. The formula for the placement of the fulcrum is f1 · d1 = f2 · d2 where f1 = force applied at one end, d1 = distance of the first force from the fulcrum, f2 = force applied at the other end, and d2 = distance of the second force from the fulcrum. Rewrite this formula to solve for d1. Show all steps in your work. F1 * D1 = F2 * D2 F1 * D1/F1 = F2 * D2/F1 D1 = F2 * D2/F1

OpenStudy (anonymous):

@Hero @mathslover @dan815 @hartnn @eliassaab @amistre64 @abb0t

OpenStudy (anonymous):

It's correct ^^

OpenStudy (anonymous):

ok how bout this one Ohm’s Law states that the voltage across a wire is given by the formula V = IR where V = voltage, I = amperage, and R = resistance. Solve V = IR for R, and use your formula to determine the resistance needed to produce 110 volts at 10 amps. Show all steps in your work. V = IR V/I = RI/I R = V/I 110 = 10R 110/10 = 10R/10 R = 110/10 11 = R

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok thx